OpenStudy (anonymous):

How do I find the ordered pair that is the solution to these equations. 3x - 2y = 8 2x + 5y = -1

OpenStudy (the_fizicx99):

You could plot them into a graphing calculator or solve them :p I'd use desmos, its a lot faster

OpenStudy (the_fizicx99):

Where they intersect is the solution.

OpenStudy (the_fizicx99):

3(2) - 2(-1) = 6 + 2 = 8 8 = 8 <-- so this works for the first equation 2(2) + 5(-1) = 4 - 5 = -1 -1 = -1 <-- also works.

OpenStudy (anonymous):

thanks (:

OpenStudy (the_fizicx99):

yw